# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v3

@@ 5 Experimental Results and Analysis
Table 1 presents main results on Paper2CodeBench, from which we observe that PaperCoder consistently outperforms all baselines on both the reference-based and reference-free evaluation settings by substantial margins. We hypothesize that this performance gap of the proposed PaperCoder stems from its top-down behavior, analyzing full papers thoughtfully before starting code generation, unlike prior approaches that typically follow the bottom-up strategy, which begins with and expands short requiremental descriptions (via role-playing or Standardized Operating Procedures). In other words, the top-down approach, operationalized within PaperCoder through the sequential process of planning, analysis, and coding, is effective in handling long-form scientific documents, which are often loosely structured from a software engineering perspective (See Table 6 for more analyses). Also, when compared to the non-comparable Oracle setting (which performs evaluations on the author-released code repositories), PaperCoder achieves performance that is on par, with only marginal differences without a statistically significant gap, demonstrating the effectiveness of our approach in faithfully implementing code whose quality is closer to the implementation by authors.
@@
Recall that the reference-free evaluation protocol is designed for cases where the ground-truth repository is not available, and to investigate whether it works as a reliable proxy for the reference-based evaluation protocol, we measure their rank correlation on all samples from Paper2CodeBench. Then, as shown in Figure 3, there is a strong positive correlation between the reference-based and reference-free evaluations, achieving a Pearson correlation coefficient of $r=0.79$. This result supports that the reference-free evaluation (despite the absence of ground-truth repositories) can serve as a reliable proxy for the reference-based evaluation, ultimately functioning as a standalone metric to assess the code quality.
@@
In addition to automatic model-based evaluations, we conduct human evaluations and report the results in Table 2. From this, we confirm that PaperCoder achieves the best ranking among all compared methods, alongside the best scores with model-based evaluations, which reaffirms its effectiveness. Moreover, to ensure whether the model-based evaluations are a reasonable proxy to judge the implementation quality for papers, we measure their correlations with human evaluation scores. As shown in Table 5, we observe strong rank correlations across both reference-based and reference-free settings (where GPT-4o and o3-mini-high are used as evaluation models), which suggests that model-based evaluation can reliably approximate human judgment. Also, based on this result, we use o3-mini-high as the default evaluation model in our experiments due to its highest correlation. Lastly, we ensure the quality and reliability of human evaluations by measuring the inter-annotator agreement score based on Cohen’s kappa coefficient, which exhibits a high score of 0.79, indicating strong consistency.
@@
Extending the model variations results on PaperBench Code-Dev, we conduct an auxiliary analysis with DS-Coder (DeepSeek-Coder-V2-Lite-Instruct [8]), Qwen-Coder (Qwen2.5-Coder-7B-Instruct [15]), DS-Distill-Qwen (DeepSeek-R1-Distill-Qwen-14B [9]), and o3-mini-high (the high reasoning-effort variant of o3-mini) on Paper2CodeBench. As summarized in Table 4, the proprietary model (o3-mini-high) consistently outperforms all other backbones across all evaluation settings. Among other open-source models, DS-Distill-Qwen performs the best, followed by Qwen-Coder and DS-Coder. These results suggest the importance of selecting a capable backbone to instantiate PaperCoder, particularly one with strong reasoning capabilities. Also, based on this result, we primarily use o3-mini-high as the basis backbone model for all experiments.
Fine-Grained Analysis of Generated Repositories
To more thoroughly evaluate the quality and practical utility of the generated code, we conduct a set of fine-grained human analyses, according to two perspectives: its usability in reproducing the original work and its component-wise implementation quality. Specifically, we first ask the human annotators whether the top-ranked repository generated by PaperCoder would make it easier to reproduce the methods and experiments compared to starting from scratch. Then, a majority of participants (92%) agree that it would facilitate reproduction, reinforcing the real-world usefulness of our approach. Also, we conduct a component-level implementation analysis to assess which parts of the papers are most effectively translated into code, by asking human annotators to identify the key component-specific elements for Data Processing, Method, and Evaluation, and then measure how many of these elements are actually implemented in the generated code. As shown in Figure 5, we find that the sections of Method, Evaluation, and Data Processing achieve the highest coverage scores, in that order, with each exceeding 50%. Notably, among the errors observed, many of them stem from missing or incorrect implementation details, often rooted in uncertainties about the structure and format of the data, as well as how the data is loaded, which are often under-specified in the papers. Lastly, to investigate why human annotators prefer PaperCoder over its baselines and ablated variants (with 22 out of 25 selecting the repositories from PaperCoder), we ask them to provide the reasons for their choices, and the majority of which are completeness, clean structure, and faithfulness to the original papers, summarized in Table 10.
Ablation Studies
To see how much each component of PaperCoder contributes to the performance gain, we conduct ablation studies on the subset of Paper2CodeBench (composed of ICML papers). Specifically, we start with the base method that uses only the full paper (called Paper) and incrementally add components in the order they are executed (such as overall plan, architecture design, logic design, configuration generation, and final analysis), whose results are reported in Table 6. From this, we observe that the performance steadily improves as additional components are incorporated. Also, we observe a performance drop when the architecture design module is added; however, while this might seem surprising at first, it is in fact expected: architecture design alone does not specify the execution or implementation order of files, which leads to confusion during the code generation stage. However, this issue is effectively addressed once the subsequent logic design module explicitly defines file dependencies and establishes a clear generation order, leading to a substantial performance boost. Overall, integrating all modules in the pipeline yields the highest performance, confirming the effectiveness of our fully structured, multi-stage pipeline, with each module playing a valuable role.
Correlation on Paper Type
To see whether the acceptance category (or presentation format) of papers correlates with the quality of their corresponding implementations by PaperCoder, we analyze it by separating papers into oral/spotlight and poster categories on Paper2CodeBench (which includes 14 oral or spotlight papers and 76 poster papers). As shown in Figure 5, scores are slightly higher for oral/spotlight papers on model-based evaluations with GPT-4o and o3-mini, suggesting that papers with higher recognition might reflect clearer writing, probably leading to faithful code generation.
Figure 4: Detailed, fine-grained analyses on code by PaperCoder.
Table 6: Ablation results on the subset of Paper2CodeBench (ICML papers) with scores and standard deviations.
@@
Figure 5: Model-based evaluation results by presentation type of papers (poster vs. oral/spotlight).
@@
It is worth noting that making the repository-level code executable and fully reproducible in one go is extremely challenging (even for human developers), as demonstrated by prior findings from Starace et al. [38]. Also, our goal is to provide a strong and faithful starting point that meaningfully aids reproduction efforts (supported by our results), rather than aiming for perfect reproduction out of the box. Nevertheless, to assess how close our generated repositories are to being directly executable, we perform manual execution evaluations on five papers. Specifically, when execution errors occur, we manually debug and refine the code and adapt the input data as needed to enable successful runs. We then find that, on average, only 0.81% of the code lines require minor modification, such as updating deprecated API calls or correcting data type mismatches, for successful execution (See Examples in Figures 6 and 7, alongside the detailed statistics in Table 9), which highlights that our generated repositories are near-executable with minimal human intervention.
@@ A.1 Implementation Details
All experiments are conducted using the o3-mini with high reasoning effort version (o3-mini-high) as the default backbone, released on January 31, 2025. To collect paper metadata and content, we use openreview_scraper [^4] https://github.com/pranftw/openreview_scraper with the OpenReview API [^5] https://docs.openreview.net/reference/api-v2 and the Semantic Scholar API [^6] https://www.semanticscholar.org/product/api. For document processing, we convert papers into structured JSON format using the s2orc-doc2json library [23] [^7] https://github.com/allenai/s2orc-doc2json. Notably, with o3-mini-high to generate repositories for 90 papers, the total API cost of PaperCoder amounts to \$76.65, resulting in an average cost of approximately \$0.90 per paper.
@@ A.3 Reference-Based Evaluation
In the reference-based evaluation setup, the repository may exceed the context length of (even frontier) LLMs. Following Starace et al. [38], when this occurs, we prompt the model to select the most relevant files for evaluation. The selected subset is then used as the reference for scoring. We use the gpt-4o-2024-11-20 as the evaluation model.
@@ A.4 PaperBench Code-Dev Evaluation
While PaperCoder is designed to generate only the source code, the PaperBench Code-Dev benchmark used for evaluation requires an additional script file called reproduce.sh. To meet this requirement, we further prompt the coding agent to generate it for this benchmark and evaluate the code with it.
@@ B.1 Code Availability
Table 8: Average Replication Scores (%) on PaperBench Code-Dev. For all OpenAI models, the reasoning effort is set to high, and we take results for BasicAgent and IterativeAgent from the PaperBench paper [38]. For PaperCoder, we report the average and standard deviation over three runs, except for o1 and o3 due to costs.
@@ Appendix C Limitations and Future Work
Appendix D Broader Impacts and Ethics Statement
Table 9: Executability results across generated repositories. We manually select five papers and generate corresponding repositories using PaperCoder. For each repository, we report the number of lines modified during debugging, the total number of lines, and the percentage of modified lines.
@@ Appendix D Broader Impacts and Ethics Statement
Table 10: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators select the repositories generated by our framework as their top choice into six (described in the first row).
@@
Figure 6: Error case from Self-Instruct. The left shows the original code generated by PaperCoder using o3-mini-high, and the right shows the version manually edited by the authors to correct the error. In this example, an outdated API call is updated to its latest version. In the original version, lines 2, 12, 14, 15, and 29 were removed; in the edited version, lines 2, 9, 10, 14, 16, 17, and 22 were added.
@@
Figure 7: Error case from CoLoR. The left shows the original code generated by PaperCoder using o3-mini-high, and the right shows the manually edited version by the authors. In this example, a numeric value is cast correctly and a required argument is added to enable execution. Lines 2 and 5 are modified in the corrected version.
Appendix E Prompts

### v4

@@ 4.2 Experimental Results and Analysis
Table 1 presents main results on Paper2CodeBench, in which PaperCoder consistently outperforms all baselines. We hypothesize that this performance gap stems from its top-down behavior, analyzing full papers thoughtfully before generation, unlike prior approaches that typically follow a bottom-up strategy, which begins with and expands short requiremental descriptions (via role-playing or SOP). In other words, the top-down approach, operationalized through the sequence of planning, analysis, and coding, is effective in handling long-form scientific documents, which are often loosely structured from a software engineering perspective. Also, when compared to the non-comparable Oracle setting (which performs evaluations on the author-released repositories), PaperCoder achieves performance that is on par, without statistically significant differences, demonstrating its effectiveness in faithfully implementing code whose quality is closer to the implementation by authors.
@@
Recall that the reference-free evaluation protocol is designed for cases where the ground-truth repository is not available, and to investigate whether it works as a reliable proxy for the reference-based evaluation protocol, we measure their rank correlation on all samples from Paper2CodeBench. Then, as shown in Figure 3, there is a strong positive correlation between them, achieving a Pearson correlation coefficient of $r=0.79$. This result supports that the reference-free evaluation can serve as a reliable proxy for the reference-based evaluation, ultimately functioning as a standalone metric to assess the code quality.
@@
In addition to automatic evaluations, we conduct human evaluations and report the results in Table 2. From this, we confirm that PaperCoder achieves the best ranking, consistent with model-based evaluations, which reaffirms its effectiveness. Also, to ensure whether the model-based evaluations are a reasonable proxy to judge the implementation quality, we measure their correlations with human evaluation scores. As shown in Table 5, we observe strong rank correlations across both reference-based and reference-free settings, which suggests that model-based evaluation can reliably approximate human judgment. Also, based on this result, we use o3-mini-high as the default evaluation model. Lastly, we ensure the quality and reliability of human evaluations by measuring the inter-annotator agreement based on Cohen’s kappa coefficient, which exhibits a high score of 0.79, indicating strong consistency.
@@
Extending the model variations results on PaperBench Code-Dev, we conduct an auxiliary analysis with DS-Coder (DeepSeek-Coder-V2-Lite-Instruct; DeepSeek-AI et al., 2024), Qwen-Coder (Qwen2.5-Coder-7B-Instruct; Hui et al., 2024), DS-Distill-Qwen (DeepSeek-R1-Distill-Qwen-14B; DeepSeek-AI et al., 2025), and o3-mini-high (the high reasoning-effort variant of o3-mini) on Paper2CodeBench. As summarized in Table 4, the proprietary model (o3-mini-high) consistently outperforms all other backbones across all evaluation settings. Among other open-source models, DS-Distill-Qwen performs the best, followed by Qwen-Coder and DS-Coder. These results suggest the importance of selecting a capable backbone to instantiate PaperCoder, particularly one with strong reasoning capabilities. Also, based on this, we primarily use o3-mini-high as the basis.
Table 6: Ablation results on the subset of Paper2CodeBench with scores and standard deviations.
@@
Table 7: Results of the PaperCoder and PaperCoder with Self-Refine, under the reference-based evaluation protocol.
PaperCoder w/ Self-Refine
Overall Plan 4.67 4.87 (+0.20)
Arch. Design 3.20 3.96 (+0.76)
Logic Design 4.09 4.38 (+0.29)
Config File 2.93 3.93 (+1.00)
Analysis 4.18 4.32 (+0.14)
Code 3.39 3.89 (+0.50)
Figure 4: Model-based evaluation results by paper presentation types.
Ablation Studies
To see how much each component of PaperCoder contributes to the performance gain, we conduct ablation studies on the subset of Paper2CodeBench (composed of ICML papers). Specifically, we start with the method that uses only the full paper and incrementally add components in the order they are executed (such as overall plan, architecture design, logic design, configuration generation, and final analysis), reported in Table 6. We then observe that the performance steadily improves as additional components are incorporated. Meanwhile, a performance drop occurs when the architecture design module is added; however, while this might seem surprising at first, it is in fact expected: architecture design alone does not specify the execution or implementation order of files, which leads to confusion during the code generation stage. However, this issue is addressed once the subsequent logic design module explicitly defines file dependencies and establishes a clear generation order. Overall, integrating all modules in the pipeline yields the highest performance, confirming the effectiveness of our fully structured, multi-stage pipeline with various modules proposed.
Experiment with Refinement
We confirm in Table 6 that the planning and analysis stages play a pivotal role in guiding subsequent analysis and coding, and we further test whether refining earlier outputs can improve downstream performance. Specifically, we augment the planning and analysis phases with verification-and-refinement steps (See Figures 19 to 28 for prompts), following Self-Refine (Madaan et al., 2023), and evaluate a total of 30 papers subsampled from Paper2CodeBench (10 from each conference). As shown in Table 7, refinement of planning and analysis improves their own outputs but also leads to measurable gains in the subsequent stages, reducing downstream errors.
Correlation on Paper Type
To see whether the acceptance category (or presentation format) of papers correlates with the quality of their corresponding implementations by PaperCoder, we analyze it by separating papers into oral/spotlight and poster categories on Paper2CodeBench (which includes 14 oral or spotlight papers and 76 poster papers). As shown in Figure 4, scores are slightly higher for oral/spotlight papers on model-based evaluations with GPT-4o and o3-mini, suggesting that papers with higher recognition might reflect clearer writing, probably leading to faithful code generation. For further analysis on how the completeness of papers impacts the results, please refer to Table 11.
Fine-Grained Analysis of Generated Repositories
To more thoroughly evaluate the quality and practical utility of the generated code, we conduct a set of fine-grained human analyses according to its usability for reproduction and its component-wise implementation quality. Specifically, we ask annotators whether the top-ranked repository from PaperCoder would make reproducing the original work easier than starting from scratch, and 92% agree, highlighting its practical value. Also, we conduct a component-level analysis to assess which parts of the papers are most effectively translated into code, by asking human annotators to identify key elements for Data Processing, Method, and Evaluation, then measure how many are actually implemented. As shown in Figure 6, the coverage reaches 80% for Method and 79% for Evaluation. Notably, among the errors observed, many of them originate from the Data Processing stage, where papers often under-specify details about data formats, preprocessing steps, or loading procedures. Lastly, to investigate why human annotators prefer PaperCoder over its baselines and ablated variants (with 22 out of 25 selecting the repositories from PaperCoder), we ask them to provide the reasons for their choices, and the majority of which are completeness, clean structure, and faithfulness to the original papers, summarized in Table 14.
4.3 Additional Analysis on Reproduction from Implemented Code Repository
While our focus is on generating faithful implementations that can aid research, we further examine whether these implementations can fully reproduce the original experimental results end-to-end.
@@ 4.3 Additional Analysis on Reproduction from Implemented Code Repository
It is worth noting that making the repository-level code executable and fully reproducible in one go is extremely challenging (even for humans), as demonstrated by Starace et al. (2025). Also, our goal is to provide a faithful starting point that meaningfully aids reproduction efforts (Figure 6), rather than aiming for perfect reproduction. Nevertheless, to assess how close our generated repositories are to being directly executable, we perform manual execution evaluations on five papers. Specifically, when execution fails, we manually debug and refine the code and adapt the input data as needed to enable successful runs. We then find that, on average, only 0.81% of the code lines require minor modification, such as updating deprecated API or correcting data type mismatches, for successful execution (See Examples in Figures 7 and 8 with statistics in Table 13), which highlights that our generated repositories are near-executable with minimal human intervention.
Figure 5: Fine-grained analyses on code by PaperCoder.
Table 8: Replication scores on 10 papers from PaperBench, including execution and result match.
Model Score (%)
BasicAgent 2.60
IterativeAgent 11.22
PaperCoder 28.46
Figure 6: Results on the author-written rubric for papers from Paper2CodeBench (human evaluated), with gains in parentheses.
Analysis on Reproducibility
An equally important, though not our primary focus, question is whether the generated repositories can reproduce the results intended by the original authors. To examine this, we sample 10 papers from PaperBench and another 10 from the human evaluation set of Paper2CodeBench. Also, we automatically invoke LLM-assisted debugging (only when execution errors occur), where the model was provided with error messages, source code, and relevant training data (if needed) to resolve issues. First, for PaperBench, we use the full rubric provided, including the aspects of result match as well as code development and execution, with o3-mini serving as the judge. Then, as shown in Table 8, PaperCoder achieves the highest score. Also, for Paper2CodeBench, we adopt the rubric defined by the paper authors, covering Data Processing, Method, and Evaluation, with o4-mini as the judge, and as shown in Figure 6, PaperCoder outperforms all baselines regardless of whether debugging is used. These results show that its repositories are not only executable with minimal (and automatically debuggable) intervention but also more faithfully reproduce the papers.
Case Study
We further conduct a manual case study on five repositories, where annotators check whether the returned outputs match the reported results. As described in Table 15 with Appendix A.5, four reproduce results (at least partially), while one fails due to issues in loss function design.
@@ 5 Conclusion
Ethics Statement
Reproducibility Statement
We attach the code to reproduce our work in the supplementary materials. Detailed instructions for running the experiments are included in the accompanying README files, and furthermore, all necessary details to reproduce our experiments are described in Section 4.1 and in Appendix A.1.
Appendix
@@ A.1 Implementation Details
All experiments are conducted using o3-mini with high reasoning effort version (o3-mini-high) as the default backbone, released on January 31, 2025. To collect paper metadata and content, we use openreview_scraper [^4] https://github.com/pranftw/openreview_scraper with the OpenReview API [^5] https://docs.openreview.net/reference/api-v2 and Semantic Scholar API [^6] https://www.semanticscholar.org/product/api. For document processing, we convert papers into structured JSON format using the s2orc-doc2json library (Lo et al., 2020) [^7] https://github.com/allenai/s2orc-doc2json. Notably, with o3-mini-high to generate repositories for 90 papers, the total API cost of PaperCoder amounts to \$76.65, resulting in an average cost of approximately \$0.90 per paper.
@@ A.3 Reference-Based Evaluation
In the reference-based evaluation setup, the repository may exceed the context length of (even frontier) LLMs. Following Starace et al. (2025), when this occurs, we prompt the model to select the most relevant files for evaluation. The selected subset is then used as the reference for scoring. We use the gpt-4o-2024-11-20 as the evaluation model.
@@ A.4 PaperBench Code-Dev Evaluation
While PaperCoder is designed to generate only the source code, the PaperBench Code-Dev benchmark used for evaluation requires an additional script file called reproduce.sh. To meet this requirement, we further prompt the coding agent to generate it and evaluate the code with it.
A.5 Additional Details on Execution and Reproducibility Experiments
To assist the reproduction of repositories from PaperCoder, we perform LLM-assisted automatic debugging. Specifically, we primarily use o4-mini for debugging, with GPT-5 used as a fallback when identical errors persist. Furthermore, all executions are performed in a Docker environment with an NVIDIA GeForce RTX 2080 GPU, and for experiments requiring larger memory, an NVIDIA RTX A6000. Lastly, due to hardware constraints, we adjust certain hyperparameters (e.g., batch size or learning rate), and in rare cases, subsampled the training data to enable successful execution. We provide the prompts in Figure 18, and statistics on the number of modified lines in Table 16.
@@ B.1 Code Availability
Table 10: Average Replication Scores (%) on PaperBench Code-Dev. For all OpenAI models, the reasoning effort is set to high, and we take results for BasicAgent and IterativeAgent from Starace et al. (2025). For PaperCoder, we report the average and standard deviation over three runs, except for o1 and o3 due to costs.
@@ B.2 PaperBench Code-Dev Results
B.3 Impact of Paper Content on Code Generation
Table 11: Comparison of reference-based average scores between the full paper content and the paper content without the methodology section on the subsampled Paper2CodeBench. Values in parentheses indicate the standard deviation.
Full (Original) w/o Methodology
Ref-based
Average Score
4.26 (0.28) 3.75 (0.55)
To examine the extent to which the clarity and specificity of the paper content influence code generation quality, we remove the Methodology section from each paper and use PaperCoder to generate the corresponding code repository. Specifically, this experiment is conducted with 30 papers (10 from each conference) in Paper2CodeBench, with o3-mini-high as the backbone LLM. As shown in Table 11, the average score drops from 4.26 to 3.75 without the Methodology section, indicating that when detailed specifications are absent, the generated code quality degrades substantially, which supports the importance of precise and explicit descriptions for faithful paper-to-code generation, as well as for human readers seeking to understand and reproduce the work.
B.4 Most Common Types of Errors and Failure Modes
Table 12: Categories of error types observed when running Paper2CodeBench. Categories are analyzed using o4-mini-high, and Count indicates the number of papers belonging to each category.
Category Count Category Count
MissingDependency 23 ConfigurationErr1or 5
ImportError 14 SyntaxError 4
ModuleNotFoundError 14 Success 4
ValueError 6 OSError 4
FileNotFoundError 6 TypeError 2
RuntimeError 6 AttributeError 2
To analyze failure cases, we execute the generated repositories on Paper2CodeBench (without debugging) and inspect the resulting errors. We note that each error is automatically categorized by prompting o4-mini-high with the raw error message and mapping its response to a canonical taxonomy. As summarized in Table 12, the most frequent causes are MissingDependency, ImportError, and ModuleNotFoundError, in that order. This pattern suggests that environment and packaging issues dominate over algorithmic or logic errors in practice.
B.5 Analysis of Performance Across Paper Categories
Examining performance across different paper categories helps reveal where code generation is easier or more challenging. To achieve this, we categorize 90 papers in Paper2CodeBench using o4-mini-high, and then report the average reference-based scores per category in Figure 9. First, we observe that the scores range from 3.38 to 4.21 (a maximum gap of about 0.83). Specifically, theory and interpretability/explainability achieve the highest scores (4.21 and 3.97), while reinforcement learning/control and dataset-focused papers yield the lowest (3.38 each). These results suggest that there are measurable variations across different categories of papers when implementing them with PaperCoder, with some types of papers being easier for PaperCoder to implement than others.
@@ Appendix C Limitations and Future Work
Table 13: Executability analysis results on the repositories: we sample five papers and generate corresponding repositories using PaperCoder. For each repository, we report the number of lines modified during debugging, the total number of lines, and the percentage of modified lines.
@@
Table 14: Qualitative analysis of top-ranked repositories. We categorize the reasons why human annotators select the repositories generated by our PaperCoder framework as their top choice into six (described in the first row).
@@
Table 15: Analysis of the results from the reproducibility case study.
Repo Name
Analysis of Reproducibility
CoLoR
Execution was successful, but the ORPO loss was likely mis-specified, causing the compression model to fail in training as intended. This issue stems from the overly simplified description of the loss function in the paper.
cognitive-behaviors
Successfully reproduced SFT and RL training processes but encountered a minor error in parsing model responses during evaluation.
RADA
Implementation closely matched the paper, but missing details prevented full reproduction of the reported results, leading to identical samples.
Self-Instruct
Executed smoothly and accurately reflected the procedure described in the paper.
G-EVAL
Implemented only the Coherence metric, though the original paper included Coherence, Consistency, Fluency, and Relevance. The Coherence implementation was faithful and correct.
Table 16: Comparison of the total lines, modified lines, and percentages when applying automatic debugging on 10 papers from Paper2CodeBench used for human evaluation.
Abstract Paper MetaGPT ChatDEV PaperCoder
Modified lines (*.py, *.sh, *.yaml) 30 705 226 275 780
Total lines 3517 3047 8225 4185 16189
Percentage 0.85 23.14 2.75 6.57 4.82
@@
Figure 7: Case study on the reproduction of the Self-Instruct paper. The left shows the code generated by PaperCoder using o3-mini-high, and the right shows the version manually edited by the authors to correct the error. In this example, an outdated API call is updated to its latest version. In the initial version, lines 2, 12, 14, 15, and 29 are removed; in the edited version, lines 2, 9, 10, 14, 16, 17, and 22 are added.
@@
Figure 8: Case study on the reproduction of the CoLoR paper. The left shows the code generated by PaperCoder using o3-mini-high, and the right shows the manually edited version by the authors. In this example, a numeric value is cast correctly, and a required argument is added to enable execution. Lines 2 and 5 are modified.
Figure 9: Average scores (measured by reference-based evaluation) per category on Paper2CodeBench. The numbers to the right of each bar indicate the average score, along with the number of papers in parentheses. Bar transparency is proportional to the count, highlighting categories with more or fewer papers.
Appendix D Prompts